OpenStudy (anonymous):

A truck is to be driven 200 km along a level highway at x km per hour. Petrol costs 8 cents per litre and is used at the rate of \[25+\frac{ x^2 }{ 112 }\] litres per hour. The driver receives 2 dollars per hour. What is the most economical speed (to the nearest km per hour and the cost of the trip- i) if there is no speed limit, ii) if the speed must not exceed 60km per hour?

OpenStudy (anonymous):

Topic: Calculus Sub Topic: Maxima and Minima.

OpenStudy (anonymous):

Is it 74.83 km/h?

OpenStudy (anonymous):

Yes!!

OpenStudy (anonymous):

\[C(x) = 0.08(25 + \frac{ x^{2} }{ 112 }) * \frac{ 200 }{ x } + 2(\frac{ 200 }{ x })\] This is the equation I came up with for Cost.

OpenStudy (anonymous):

75km/hour

OpenStudy (anonymous):

How did you do the first part?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

gallon?

OpenStudy (anonymous):

American?

OpenStudy (anonymous):

let me rewrite that , (i used american units and accidentally wrote it in the wrong place)

OpenStudy (anonymous):

The first thing we know is that the rate of petrol is $0.08/litre And the second thing we know is that the petrol is used at a rate of 25 + x^2/112 litre/h so \[0.08 \frac{ dollar}{ litre } * (25 + \frac{ x^{2} }{ 112 }) \frac{ litre }{ hour }\] Gives us the units dollars/hour

OpenStudy (anonymous):

The only thing we need to know is how long the trip takes. We know speed = distance / time speed = x distance = 200 km time = t Therefore time, t = 200/x If we multiply what we have so far by (200/x) (which is in hours), then the hours unit cancels leaving an expression in dollars

OpenStudy (anonymous):

So far, I'm up to: \[0.08*(25+\frac{ x^2 }{ 112 })\]

OpenStudy (anonymous):

\[0.08 \frac{ dollar }{ litre } * (25 + \frac{ x^{2} }{ 112 }) \frac{ litre }{ hour } * \frac{ 200 }{ x } hours\]

OpenStudy (anonymous):

So you get that, and do you then add: \[\frac{ 400 }{ x }\] for the money he receives per hour?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay. Now I'm getting somewhere.

OpenStudy (anonymous):

So do we differentiate that with respect to x?

OpenStudy (anonymous):

Yes because C(x) is the cost evaluated at x. So we need to minimize cost by differentiating with respect to x.

OpenStudy (anonymous):

Thanks for the assistance @LogicalApple . Appreciate it.

OpenStudy (anonymous):

Not a problem

OpenStudy (anonymous):

For the second part of the question, keep in mind that the point we evaluated in the first part was a minimum. Due to the nature of the hyperbola, you can consider it an absolute minimum whenever x > 0 This should help you answer the second part of the question.

OpenStudy (anonymous):

Yeah I already finished the question.

OpenStudy (anonymous):

Oh, nice

OpenStudy (anonymous):

The hard part was what you did for me.

OpenStudy (anonymous):

I did it on paper and I was like "hm.. I can't think in litres/gallon"

OpenStudy (anonymous):

Ah, I did it again lol I mean litres/hour!

OpenStudy (anonymous):

hahah.

OpenStudy (anonymous):

Well I'm going to close this and put up another question. Thanks again for helping me.

OpenStudy (anonymous):

Hope you can help me again if you see fit to.

OpenStudy (anonymous):

Good luck!